# Pricing Options on Forwards in Function-Valued Affine Stochastic Volatility Models

### v1

@@ 3. Pricing options on forwards in infinite-dimensional stochastic volatility models
From now on we impose the no–arbitrage drift condition (3.6). Equivalently, in the coupled SV dynamics (2.10) or (2.4) under $\mathbb{Q}$ we choose
$\Upsilon^{\vartheta}:=-\tfrac{1}{2}\,S^{*}(\vartheta)h_{0},\qquad\vartheta\geq 0,\ \ h_{0}\in H_{w}.$ (3.7)
Indeed, with this parametrization the model drift satisfies $C_{t}-D^{1/2}Y_{t}^{1/2}\xi_{t}=D^{1/2}Y_{t}D^{1/2}\Upsilon^{\vartheta}$, and for a claim with delivery lag $\vartheta=\tau-t$ this equals $\tfrac{1}{2}D^{1/2}Y_{t}D^{1/2}S^{*}(\tau-t)h_{0}$, which is exactly (3.6). Moreover, $\vartheta\mapsto D^{1/2}S^{*}(\vartheta)h_{0}\in H_{w}$ (see [7, Lem. 6.2]), so $\Upsilon^{\vartheta}\in H_{w}$ is well defined.
@@
Proposition 3.2.
Let $T>0$, $0\leq T_{0}\leq T_{1}\leq T$, and $h_{0}\in H_{w}$. Set $\vartheta:=T_{1}-T_{0}$. Assume $(X^{\vartheta}_{t})_{t\geq 0}$ is as in (3.2), with $\Upsilon=-\tfrac{1}{2}S^{*}(\vartheta)h_{0}$, and let the forward price $F$ be defined by (3.8). Write $\mathbb{E}_{\mathbb{Q}}$ for expectation under $\mathbb{Q}$, and define $g:\mathbb{R}\to\mathbb{C}$ by
@@
Set $u:=(\nu+\mathrm{i}\lambda)\,u_{\vartheta}$. We consider the two covariance specifications below.
  1. (i)
    If the covariance process is $Y_{t}=\mathsf{Y}_{t}$ as in (2.2), and $(\phi^{\vartheta},\psi_{1}^{\vartheta},\psi_{2}^{\vartheta})$ denotes the solution to the generalized Riccati equations (2.8a)–(2.8c) on $[0,T]$ with initial data $(u,0)$, then for $0\leq t\leq T_{0}$ the time-$t$ price of the call with maturity $T_{0}$ and strike $K>0$ written on $F$ is
@@
    $\hskip 45.52458pt-\langle\psi_{2}^{\vartheta}(T_{0}-t,u,0),\mathsf{Y}_{t}\rangle_{\mathcal{H}^{\mathbb{C}}}\Big{)}\,d\lambda.$ (3.10)
  2. (ii)
    If the covariance process is $Y_{t}=\mathcal{Y}_{t}$ as in (2.2.1), and $(\eta^{\vartheta},\zeta_{1}^{\vartheta},\zeta_{2}^{\vartheta})$ denotes the solution to the Riccati equations (2.15a)–(2.15c) on $[0,T]$ with initial data $(u,0)$, then for $0\leq t\leq T_{0}$
    $\mathbb{E}_{\mathbb{Q}}\big{[}(F(T_{0},T_{1})-K)^{+}\mid\mathcal{F}_{t}\big{]}=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\!\Big{(}$ $-\eta^{\vartheta}(T_{0}-t,u,0)+\langle\zeta_{1}^{\vartheta}(T_{0}-t,u,0),X_{t}^{\vartheta}\rangle_{H_{w}^{\mathbb{C}}}$
    $\hskip 45.52458pt-\langle\zeta_{2}^{\vartheta}(T_{0}-t,u,0),\mathcal{Y}_{t}\rangle_{\mathcal{H}^{\mathbb{C}}}\Big{)}\,d\lambda.$ (3.11)
@@
Standard Fourier techniques for option pricing (see e.g., [23] or [26, Theorem 10.6]), yields
$\mathbb{E}_{\mathbb{Q}}[(F(T_{0},T_{1})-K)^{+}|\mathcal{F}_{t}]=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\mathbb{E}_{\mathbb{Q}}[\operatorname{e}^{\langle X^{\vartheta}_{T_{0}},(\nu+\mathrm{i}\lambda)u_{\vartheta}\rangle_{H_{w}^{\mathbb{C}}}}|\mathcal{F}_{t}]\,\mathop{}\!\textnormal{d}\lambda\,.$
Then the statement follows from Theorems 2.3 and 2.6. ∎
Consider a put option with strike $K>0$ and exercise time $T_{0}$, i.e., the payoff is given by $(K-F(T_{0},T_{1}))^{+}$. Similarly to the case of a call option, we obtain the expressions in the right hand side of ((ii)) and ((i)) for the price of a put option respectively, written on $(X_{t}^{\vartheta},\mathcal{Y}_{t})_{t\geq 0}$ and $(X_{t}^{\vartheta},{\sf Y}_{t})_{t\geq 0}$ with $u=(\nu+\mathrm{i}\lambda)u_{\vartheta}$, for $\nu<0,\lambda\in\mathbb{R}$.
@@ 4. Numerical analysis
In this section we study several specifications of the coupled process $(X_{t},Y_{t})_{t\geq 0}$ and compute prices of options written on forwards. Our primary goal is to verify, across models, that the option values obtained from the affine, semi-explicit formulas ((i)) and ((ii)) match those produced by Monte Carlo (MC) simulation within sampling error and show that that the affine formulas allows for much more efficient pricing. We start by describing the general settings of our model specifications.
@@ 4.1. General settings
In the implementation, we choose as a weight function $w(x)=\operatorname{e}^{\alpha x}$, for some variable $\alpha>0$. The initial forward rate curve is chosen using a Nelson-Siegel curve. We know that $H_{w}$ is isometrically isomorphic to $\mathbb{R}\times L^{2}[0,\infty)$ via the isomorphism
$T(h)=(h(0),h^{\prime}w^{1/2})\,,\qquad T^{-1}(c,f)=c+\int_{0}^{\cdot}f(y)w^{-1/2}\,\mathop{}\!\textnormal{d}y.$ (4.1)
Assume $(e_{n})_{n\in N}$ is an orthonormal basis of $\mathbb{R}\times L^{2}[0,\infty)$, then $(f_{n})=(T^{-1}e_{n})_{n\in N}$ is an orthonormal basis of $H_{w}$. We choose the orthonormal basis $(1,0)$ and $(0,g_{n})$ for $\mathbb{R}\times L^{2}[0,\infty)$, where $(g_{n})_{n\in\mathbb{N}}$ are Laguerre polynomials multiplied by a weight function $\mathrm{e}^{-x/2}$ (see Appendix A). It follows that the orthogonal basis $(f_{n})_{n\in\mathbb{N}}$ for $H_{w}$ is given by
@@
$f_{n+1}$ $=\int_{0}^{\cdot}g_{n-1}(s)w^{-1/2}(s)\,\mathop{}\!\textnormal{d}s,\quad n\geq 1\,.$
@@
We infer for $(S^{*}(t))_{t\geq 0}$ and the weight function $w(x)=\mathrm{e}^{\alpha x}$,
$S^{*}(t)h(x)$ $=h(0)+h(0)\int_{0}^{x\wedge t}w^{-1}(s)ds+\int_{t}^{x\vee t}\frac{h^{\prime}(s-t)w(s-t)}{w(s)}\mathop{}\!\textnormal{d}s$
$=h(0)+\frac{1}{\alpha}h(0)(1-\mathop{}\!\textnormal{e}^{-\alpha(x\wedge t)})+\mathop{}\!\textnormal{e}^{-\alpha t}\mathbbm{1}_{x\geq t}(h(x-t)-h(0))\,.$ (4.3)
We take the HJM drift $D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon$ in the drift of $(X_{t})_{t\geq 0}$ equal to $\frac{1}{2}D^{1/2}\mathsf{Y}_{t}D^{1/2}\mathcal{S}^{*}(\vartheta)h_{0}$ with $h_{0}\in H_{w}$. In the experiments on the stochastic volatility model with jumps, we consider two examples: the model driven by a Lévy subordinator with $B(Y_{t})=0$, and the so called Barndorff–Nielsen–Shephard (BNS) model, i.e., $B(\mathsf{Y}_{t})=(\mathbb{C}\mathsf{Y}_{t}+\mathsf{Y}_{t}\mathbb{C}^{*})$.
4.2. Validation of the MGF in the Infinite-Dimensional Wishart Stochastic Volatility Model
For the infinite-dimensional Wishart stochastic volatility model, we first validate the theoretical moment generating function (MGF) against its empirical counterpart obtained by Monte Carlo. This is the right diagnostic because, in light of (2.2.4), option prices are computed from a fixed finite-rank approximation; consistency of the MGF is therefore the primary check for pricing accuracy.
Throughout this experiment we work in a finite-rank setting: the Wishart process takes values in the cone of self-adjoint trace-class operators of rank at most $n$. We fix an arbitrary but finite $n\in\mathbb{N}$ and an orthonormal basis $(f_{k})_{k\in\mathbb{N}}$ of $H$ as introduced earlier, and project all operators onto $H_{n}:=\mathrm{span}\{f_{1},\dots,f_{n}\}$. The theoretical MGF is evaluated via the affine transform/Riccati solution on $H_{n}$, while the empirical MGF is computed from Monte Carlo samples as the sample average of $\exp\{\langle X_{T},u_{1}\rangle-\mathrm{Tr}(Y_{T}u_{2})\}$ for the same $(u_{1},u_{2})$.
@@ 4.2. Validation of the MGF in the Infinite-Dimensional Wishart Stochastic Volatility Model
$\mathop{}\!\textnormal{d}Y_{t}=nQ\,\mathop{}\!\textnormal{d}t+Y_{t}\mathbb{A}\,\mathop{}\!\textnormal{d}t+\mathbb{A}^{*}Y_{t}\,\mathop{}\!\textnormal{d}t+\sqrt{Y_{t}}\,\mathop{}\!\textnormal{d}B_{t}\sqrt{Q}+\sqrt{Q}\,\mathop{}\!\textnormal{d}B_{t}^{*}\sqrt{Y_{t}},$
@@
The Wishart process is coupled withththth the forward dynamics via the volatility-modulated SPDE given in equation (1.1).
@@
Monte Carlo simulation of the Wishart model is considerably more demanding. The main difficulty lies in computing the square root of the operator-valued volatility process $Y_{t}$ at each time step, this is analytically intractable and must be approximated numerically. To limit computational overhead, we simulate only a single time step corresponding to a short horizon of $T_{0}=1/365$. However, at such a small horizon $T_{0}$, the inverse Fourier transform used for option pricing becomes numerically unstable. Therefore, instead of the option prices, we compare the moment generating function (MGF), which is the key intermediate quantity in affine models.
In this example, we find that the MGF computed from the affine representation aligns closely with the MGF estimated via Monte Carlo simulation, thereby confirming pricing consistency. This agreement is illustrated in Figure 1, where the Monte Carlo estimates converge to the affine-model MGF.
Figure 1. Moment generation function: Monte Carlo v.s. analytical formula, with $h_{0}=1,\alpha=0.1,T_{0}=1/365,N=5$.

### v2

@@ 3. Pricing options on forwards in infinite-dimensional stochastic volatility models
From now on we impose the no–arbitrage drift condition (3.6) for a fixed delivery lag $\vartheta$. Equivalently, in the coupled SV dynamics (2.10) or (2.4) used below to price claims with delivery lag $\vartheta$ under $\mathbb{Q}$ we choose
$\Upsilon^{\vartheta}:=-\tfrac{1}{2}\,u_{\vartheta}=-\tfrac{1}{2}\,S^{*}(\vartheta)h_{0},\qquad\vartheta\geq 0,$ (3.7)
Indeed, with this parametrization the model drift satisfies $C_{t}-D^{1/2}Y_{t}^{1/2}\xi_{t}=D^{1/2}Y_{t}D^{1/2}\Upsilon^{\vartheta}$, and for a claim with delivery lag $\vartheta=\tau-t$ this equals $\tfrac{1}{2}D^{1/2}Y_{t}D^{1/2}S^{*}(\tau-t)h_{0}$, which is exactly (3.6). Moreover, $\vartheta\mapsto D^{1/2}S^{*}(\vartheta)h_{0}\in H_{w}$ (see [7, Lem. 6.2]), so $\Upsilon^{\vartheta}\in H_{w}$ is well defined. Accordingly, all pricing results below are formulated under a risk-neutral specification for a fixed delivery lag $\vartheta$. This is sufficient for pricing the corresponding family of forward options, but we do not claim a single maturity-uniform HJMM drift specification for all delivery lags simultaneously.
@@
Proposition 3.2 (Call option pricing in the pure-jump affine SV model).
Let $T>0$, $0\leq T_{0}\leq T_{1}\leq T$, and set $\vartheta:=T_{1}-T_{0}$. Assume $(X^{\vartheta}_{t})_{t\geq 0}$ is as in (3.2), with $\Upsilon=-\tfrac{1}{2}u_{\vartheta}$, and let the forward price $F$ be defined by (3.8). Write $\mathbb{E}_{\mathbb{Q}}$ for expectation under $\mathbb{Q}$, and define $g:\mathbb{R}\to\mathbb{C}$ by
@@
Set $u:=(\nu+\mathrm{i}\lambda)\,u_{\vartheta}$. Assume the covariance process is $Y_{t}=\mathsf{Y}_{t}$ as in (2.2), and let $(\sf P,\sf q_{1},\sf q_{2})$ be the mild solution to the extended Riccati equations (2.7a)–(2.7c) on $[0,T]$ with initial value $(\Re(u),0)$, the existence of which is guaranteed by Proposition 5.2. Assume in addition that the Fourier integrand in (3.2) belongs to $L^{1}(\mathbb{R},\mathop{}\!\mathrm{d}\lambda)$ almost surely. Then for $0\leq t\leq T_{0}$ the time-$t$ price of the call with maturity $T_{0}$ and strike $K>0$ written on $F$ is
@@
$\hskip 45.52458pt-\langle\psi_{2}^{\vartheta}(T_{0}-t,u,0),\mathsf{Y}_{t}\rangle_{\mathcal{H}^{\mathbb{C}}}\Big)\,\mathop{}\!\mathrm{d}\lambda\,$, (3.10)
where $(\phi^{\vartheta},\psi_{1}^{\vartheta},\psi_{2}^{\vartheta})$ denotes the solution to the generalized Riccati equations (2.8a)–(2.8c) on $[0,T]$ with initial data $(u,0)$.
@@
Standard Fourier techniques for option pricing (see, e.g., [21] or [24, Theorem 10.6]) yield
$\mathbb{E}_{\mathbb{Q}}[(F(T_{0},T_{1})-K)^{+}|\mathcal{F}_{t}]=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\mathbb{E}_{\mathbb{Q}}[\mathop{}\!\mathrm{e}^{\langle X^{\vartheta}_{T_{0}},(\nu+\mathrm{i}\lambda)u_{\vartheta}\rangle_{H_{w}^{\mathbb{C}}}}|\mathcal{F}_{t}]\,\mathop{}\!\mathrm{d}\lambda\,.$
Then the statement follows from Theorem 2.5. ∎
Proposition 3.3 (Approximate call option pricing in the Wishart SV model).
Under the same hypotheses as Proposition 3.2, assume instead that the covariance process is $Y_{t}=\mathcal{Y}_{t}$ as in (2.2.1), and let $(P_{n},q_{1,n},q_{2,n})$ be the mild solution to the finite-rank Riccati equations (2.14a)–(2.14c) on $[0,T]$ with initial value $(\Re(u),0)$, the existence of which is guaranteed by Proposition 5.6. Assume in addition that the Fourier integrand in (3.3) belongs to $L^{1}(\mathbb{R},\mathop{}\!\mathrm{d}\lambda)$ almost surely. Then, for $0\leq t\leq T_{0}$, the finite-rank approximation of order $n$ to the call price is
$\mathbb{E}_{\mathbb{Q}}\big[(F(T_{0},T_{1})-K)^{+}\mid\mathcal{F}_{t}\big]\approx\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big($ $-\eta_{n}^{\vartheta}(T_{0}-t,u,0)+\langle\zeta_{1}^{\vartheta}(T_{0}-t,u,0),X_{t}^{\vartheta}\rangle_{H_{w}^{\mathbb{C}}}$
$\hskip 45.52458pt-\operatorname{Tr}\big(\mathcal{Y}_{t}\iota_{n}\zeta_{2,n}^{\vartheta}(T_{0}-t,u,0)\iota_{n}^{*}\big)\Big)\,\mathop{}\!\mathrm{d}\lambda\,$, (3.11)
where $(\eta^{\vartheta}_{n},\zeta_{1}^{\vartheta},\zeta_{2,n}^{\vartheta})$ denotes the solution to the Riccati equations (2.15a)–(2.15c) on $[0,T]$ with initial data $(u,0)$. The approximation arises from the finite-rank projection (2.2.4) of the Wishart process.
Proof.
The Fourier decomposition of the call payoff is identical to the proof of Proposition 3.2. Substituting the finite-rank approximation (2.2.4) for the conditional Laplace transform and applying Theorem 2.8 yields the result. ∎
Consider a put option with strike $K>0$ and exercise time $T_{0}$, i.e., the payoff is given by $(K-F(T_{0},T_{1}))^{+}$. Similarly to the case of a call option, the expressions in (3.2) and (3.3) yield the price of the corresponding put option on $(X_{t}^{\vartheta},{\sf Y}_{t})_{t\geq 0}$ and $(X_{t}^{\vartheta},\mathcal{Y}_{t})_{t\geq 0}$, respectively, by taking $u=(\nu+\mathrm{i}\lambda)u_{\vartheta}$ with $\nu<0$ and $\lambda\in\mathbb{R}$, provided the corresponding transform exists and the Fourier integrand is integrable. We use these damping conditions only as sufficient assumptions for Fourier inversion and do not attempt to characterize a maximal admissible strip in full generality.
@@ 4. Numerical analysis
In this section we study several specifications of the coupled process $(X_{t},Y_{t})_{t\geq 0}$ and compute prices of options written on forwards. Our primary goal is to verify, across models, that the pure-jump option values from the affine, semi-explicit formulas (3.2) and the Wishart transform quantities underlying (3.3) are consistent with direct or conditional Monte Carlo benchmarks within sampling error, and to compare their computational cost. We start by describing the general settings of our model specifications.
@@ 4.1. General settings
In the implementation, we choose as a weight function $w(x)=\mathop{}\!\mathrm{e}^{\alpha x}$, for some variable $\alpha>0$. The initial forward rate curve is chosen using a Nelson-Siegel curve. We know that $H_{w}$ is isometrically isomorphic to $\mathbb{R}\times L^{2}[0,\infty)$ via the isomorphism
$T(h)=(h(0),h^{\prime}w^{1/2})\,,\qquad T^{-1}(c,f)=c+\int_{0}^{\cdot}f(y)w^{-1/2}\,\mathop{}\!\mathrm{d}y.$ (4.1)
Assume $(e_{n})_{n\in N}$ is an orthonormal basis of $\mathbb{R}\times L^{2}[0,\infty)$, then $(f_{n})=(T^{-1}e_{n})_{n\in N}$ is an orthonormal basis of $H_{w}$. We choose the orthonormal basis $(1,0)$ and $(0,g_{n})$ for $\mathbb{R}\times L^{2}[0,\infty)$, where $(g_{n})_{n\in\mathbb{N}}$ are Laguerre polynomials multiplied by a weight function $\mathop{}\!\mathrm{e}^{-x/2}$ (see Appendix A). It follows that the orthogonal basis $(f_{n})_{n\in\mathbb{N}}$ for $H_{w}$ is given by
@@
$f_{n+1}$ $=\int_{0}^{\cdot}g_{n-1}(s)w^{-1/2}(s)\,\mathop{}\!\mathrm{d}s,\quad n\geq 1\,.$
@@
We infer for $(S^{*}(t))_{t\geq 0}$ and the weight function $w(x)=\mathop{}\!\mathrm{e}^{\alpha x}$,
$S^{*}(t)h(x)$ $=h(0)+h(0)\int_{0}^{x\wedge t}w^{-1}(s)\,\mathop{}\!\mathrm{d}s+\int_{t}^{x\vee t}\frac{h^{\prime}(s-t)w(s-t)}{w(s)}\mathop{}\!\mathrm{d}s$
$=h(0)+\frac{1}{\alpha}h(0)(1-\mathop{}\!\mathrm{e}^{-\alpha(x\wedge t)})+\mathop{}\!\mathrm{e}^{-\alpha t}\mathbbm{1}_{x\geq t}(h(x-t)-h(0))\,.$ (4.3)
Throughout the numerics we set $h_{0}\equiv 1$, so that $u_{\vartheta}=S^{*}(\vartheta)h_{0}$ is the evaluation functional at maturity lag $\vartheta$. Accordingly, we take
$\Upsilon^{\vartheta}=-\frac{1}{2}\mathcal{S}^{*}(\vartheta)h_{0},$
so that the HJM drift in $(X_{t})_{t\geq 0}$ is $D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon^{\vartheta}$. In the experiments on the stochastic volatility model with jumps, we consider two examples: the model driven by a Lévy subordinator with $B(Y_{t})=0$, and the so called Barndorff–Nielsen–Shephard (BNS) model, i.e., $B(\mathsf{Y}_{t})=(\mathbb{C}\mathsf{Y}_{t}+\mathsf{Y}_{t}\mathbb{C}^{*})$.
4.2. Validation of the First Forward Moment and Option Prices in the Finite-Rank Wishart Model
For the Wishart stochastic volatility model, we first validate the first forward moment against its empirical counterpart obtained by Monte Carlo. In the code this corresponds to the special transform input $u_{1}=u_{\vartheta}$ and $u_{2}=0$, hence to the quantity $\mathbb{E}\left[\exp(\langle X_{T},u_{\vartheta}\rangle_{w})\right]$. This is weaker than a full complex-MGF validation, but it checks the risk-neutral drift restriction and the finite-rank Riccati implementation in the simplest transform configuration relevant for pricing.
Throughout this experiment we work in the following setting: the coefficients $Q$, $\mathbb{A}$, $D$ and the initial condition $Y_{0}$ are all supported on $H_{n}:=\mathrm{span}\{f_{1},\dots,f_{n}\}$. In particular, for this numerical specification the finite-rank transform is exact on $H_{n}$ rather than merely asymptotic. The theoretical first moment is therefore evaluated via the matrix Riccati system on $H_{n}$, while the empirical first moment is computed from Monte Carlo samples as the sample average of $\exp\{\langle X_{T},u_{\vartheta}\rangle_{w}\}$.
@@ 4.2. Validation of the First Forward Moment and Option Prices in the Finite-Rank Wishart Model
$\mathop{}\!\mathrm{d}Y_{t}=nQ\,\mathop{}\!\mathrm{d}t+Y_{t}\mathbb{A}\,\mathop{}\!\mathrm{d}t+\mathbb{A}^{*}Y_{t}\,\mathop{}\!\mathrm{d}t+\sqrt{Y_{t}}\,\mathop{}\!\mathrm{d}B_{t}\sqrt{Q}+\sqrt{Q}\,\mathop{}\!\mathrm{d}B_{t}^{*}\sqrt{Y_{t}},$
@@
The Wishart process is coupled with the forward dynamics via the volatility-modulated SPDE given in equation (1.1).
@@
For Monte Carlo simulation we exploit the standard factor representation of the finite-dimensional Wishart diffusion. Writing the matrix-valued covariance process on $H_{n}$ as
$Y_{t}=U_{t}^{\top}U_{t},$
with $U_{t}\in\mathbb{R}^{n\times n}$ solving the Ornstein–Uhlenbeck SDE
$\mathop{}\!\mathrm{d}U_{t}=U_{t}A_{n}\,\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}W_{t}\,Q_{n}^{1/2},$
one recovers exactly the finite-dimensional Wishart dynamics on $H_{n}$. In particular, the simulated covariance matrices remain symmetric positive semidefinite at the grid points without any diagonal reduction. For the first-moment experiment we use the short horizon $T_{0}=1/365$ and compare the empirical average of $\exp\{\langle X_{T_{0}},u_{\vartheta}\rangle_{w}\}$ with the affine/Riccati value.
In this corrected implementation, the theoretical first moment aligns closely with its Monte Carlo estimate, thereby confirming the consistency of the finite-rank affine transform in this benchmark configuration. This agreement is illustrated in Figure 1.
Figure 1. First forward moment $\mathbb{E}\left[\exp(\langle X_{T_{0}},u_{\vartheta}\rangle_{w})\right]$: Monte Carlo vs. affine formula, with $h_{0}=1,\alpha=0.1,T_{0}=1/365,N=5$.
To address the option pricing validation, we employ a conditional Gaussian Monte Carlo scheme. Given a simulated path of the full finite-rank Wishart process $(Y_{s})_{0\leq s\leq t}$, the log-forward $\log F(t,t+\vartheta)=\langle X_{t},u_{\vartheta}\rangle_{w}$ is conditionally Gaussian with conditional mean
$m_{\mathrm{cond}}=\langle X_{0},u_{\vartheta+t}\rangle_{w}-\frac{1}{2}\int_{0}^{t}\big\langle D^{1/2}Y_{s}D^{1/2}u_{\vartheta},\,u_{\vartheta+t-s}\big\rangle_{w}\,\mathop{}\!\mathrm{d}s$
and conditional variance
$\sigma^{2}_{\mathrm{cond}}=\int_{0}^{t}\big\langle D^{1/2}Y_{s}D^{1/2}u_{\vartheta+t-s},\,u_{\vartheta+t-s}\big\rangle_{w}\,\mathop{}\!\mathrm{d}s.$
Hence a call option price can be computed path-wise via the Black formula. In contrast to the previous diagonal-CIR surrogate, this conditional Gaussian scheme retains all off-diagonal covariance terms of $Y_{s}$ through the full quadratic forms above. Numerically, the former structural pricing gap disappears after this correction: in our implementation the remaining difference between conditional Monte Carlo and the affine formula is of the order expected from Monte Carlo noise together with time/Fourier discretization, whereas the diagonal surrogate systematically overprices by omitting the Wishart cross-correlation terms.
Figure 2. Call option prices: conditional Gaussian Monte Carlo based on the full finite-rank Wishart process vs. affine formula, with $h_{0}=1,\alpha=0.1,t=1,k=1,N=5$.
We emphasize that the slight decline of the pricing curves in Figure 2 for large values of $\vartheta$ is not a numerical artefact. In this experiment the exercise time $t$ is fixed, and the horizontal axis varies the delivery lag $\vartheta$ in the forward contract $F(t,t+\vartheta)$. Hence the figure does not display option prices as a function of option maturity, and monotonicity in $\vartheta$ should not be expected. In the present specification, the deterministic initial forward level $F(0,t+\vartheta)$ decreases mildly with $\vartheta$, whereas the Wishart volatility loadings entering the conditional variance increase only up to a point and then essentially saturate. Consequently, for large $\vartheta$ the volatility contribution becomes nearly flat while the underlying forward level continues to decrease slightly, which leads to a small downward bend in the call prices. Any residual pointwise irregularity is due to Monte Carlo and Fourier discretization, but the overall long-end shape is a genuine feature of the model under this parameter choice.